# A Benchmark for Multi-modal Foundation Models on Low-level Vision: from Single Images to Pairs

### v1

@@ II-C Benchmark on Low-level Description Ability
In the second task of Q-Bench, we evaluate the language description ability of MLLMs on low-level information. This task is a sibling task of image captioning [7, 53, 54] that describes image content with natural language, with a specific concern on the low-level appearance of images. To evaluate this ability automatically, we first derive a golden low-level description dataset, denoted as LLDescribe (Sec. II-C1), including one long (average 58 words) golden description provided by experts for each of 499 images. With these golden text descriptions, we are able to measure the quality of output low-level descriptions from MLLMs with a single-modal GPT, under the three dimensions: completeness, preciseness, as well as relevance (Sec II-C2). The discussions of the golden descriptions and the evaluation process are as follows.
@@ II-C1 Defining Golden Low-level Descriptions for Images
For the description ability, MLLMs should accurately and completely describe low-level visual information of images. Thus, the ground truths for these MLLMs are also built within a basic principle to cover as many low-level concerns as possible, so long as they are enumerated in Sec. 3 and occur in images. The resulting golden descriptions in LLDescribe have an average duration of 58 words, notably longer than common high-level image caption datasets (11 for [54], 10 for [7]). Similar to the LLVisionQA dataset for the perception task, the 499 single images and 450 image pairs in LLDescribe dataset also include all 10 sources (as in Table I) to cover images with diverse low-level appearances. The golden descriptions on different sources of images are depicted in Fig. 4.
@@ II-C2 Evaluation with Single-modal GPT
Recent studies [55] have proved single-modal GPT [28] to be a reliable evaluation tool for pure language tasks. Via the LLDescribe dataset, we convert the multi-modality problem into a text-only setting, by matching the MLLM outputs with the golden descriptions with single-modal GPT under three dimensions: (1) Completeness. More matched information with the golden description is encouraged. (2) Preciseness. The controversial information with the golden description is punished. (3) Relevance. More proportions of MLLM outputs should be related to low-level information, instead of others. Each dimension is scored among [0,1,2]. Similar as Sec. II-B, we repeat 5 rounds for each single evaluation and collect the weighted average as the final score.
@@ II-D Benchmark on Precise Quality Assessment Ability
| Emu2-Chat (LLaMA-33B) [25] | 71.81% | 67.25% | 56.18% | 64.78% | 63.19% | 63.48% | 72.24% | 65.28% |
| Fuyu-8B (Persimmon-8B) [57] | 53.33% | 43.70% | 38.00% | 40.81% | 47.40% | 45.45% | 49.23% | 45.05% |
| BakLLava (Mistral-7B) [26] | 66.00% | 56.16% | 51.12% | 51.15% | 61.57% | 53.72% | 72.00% | 57.48% |
| SPHINX [58] | 74.18% | 68.81% | 62.07% | 63.62% | 71.76% | 66.12% | 76.33% | 68.56% |
| mPLUG-Owl2 (LLaMA-7B) [59] | 72.18% | 57.96% | 56.19% | 56.68% | 69.21% | 53.29% | 72.65% | 61.61% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 66.36% | 58.19% | 50.51% | 49.42% | 65.74% | 54.61% | 70.61% | 58.66% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 65.27% | 64.38% | 56.59% | 56.03% | 67.13% | 61.18% | 67.35% | 62.14% |
| InternLM-XComposer-VL (InternLM) [61] | 69.45% | 65.27% | 60.85% | 61.67% | 70.14% | 56.91% | 75.10% | 65.35% |
| IDEFICS-Instruct (LLaMA-7B) [62] | 56.18% | 44.69% | 44.02% | 42.80% | 54.17% | 44.74% | 56.33% | 48.70% |
| Qwen-VL (QwenLM) [63] | 63.09% | 58.19% | 56.39% | 50.58% | 62.73% | 57.89% | 73.88% | 59.40% |
| Shikra (Vicuna-7B) [64] | 65.64% | 47.35% | 49.09% | 48.83% | 59.49% | 50.00% | 64.08% | 54.65% |
| Otter-v1 (MPT-7B) [6] | 57.09% | 40.71% | 39.55% | 42.22% | 49.31% | 44.08% | 52.65% | 46.35% |
| InstructBLIP (Flan-T5-XL) [5] | 67.64% | 59.96% | 55.98% | 56.23% | 65.51% | 58.22% | 69.39% | 61.47% |
| InstructBLIP (Vicuna-7B) [5] | 71.64% | 52.65% | 43.81% | 48.64% | 62.50% | 55.59% | 64.90% | 56.72% |
| VisualGLM-6B (GLM-6B) [65] | 60.18% | 54.20% | 46.25% | 51.75% | 54.40% | 53.62% | 57.14% | 53.78% |
| mPLUG-Owl (LLaMA-7B) [52] | 66.00% | 54.87% | 44.02% | 51.36% | 55.09% | 54.28% | 65.71% | 55.38% |
| LLaMA-Adapter-V2 [66] | 66.18% | 59.29% | 52.13% | 57.39% | 56.25% | 63.16% | 64.90% | 59.46% |
| LLaVA-v1 (Vicuna-13B) [3] | 54.00% | 53.10% | 55.38% | 48.64% | 54.63% | 55.59% | 63.27% | 54.18% |
| MiniGPT-4 (Vicuna-13B) [4] | 55.82% | 50.22% | 40.37% | 42.02% | 48.38% | 51.97% | 61.22% | 49.03% |
@@
| Qwen-VL-Max (Close-Source) [63] | 75.60% | 79.43% | 66.09% | 73.39% | 74.08% | 71.0% | 76.92% | 73.63% |
| Gemini-Pro (Close-Source) [27] | 68.80% | 73.74% | 62.34% | 66.30% | 71.34% | 63.91% | 73.09% | 68.16% |
| GPT-4V (Close-Source) [28] | 76.85% | 79.17% | 67.52% | 73.53% | 76.18% | 72.83% | 76.47% | 74.51% |
@@
| Emu2-Chat (LLaMA-33B) [25] | 70.09% | 65.12% | 54.11% | 66.22% | 62.96% | 63.47% | 73.21% | 64.32% |
| Fuyu-8B (Persimmon-8B) [57] | 62.22% | 35.79% | 36.62% | 41.07% | 49.40% | 45.89% | 49.04% | 45.75% |
| BakLLava (Mistral-7B) [26] | 66.46% | 61.48% | 54.83% | 51.33% | 63.76% | 56.52% | 78.16% | 61.02% |
| SPHINX [58] | 74.45% | 65.50% | 62.13% | 59.11% | 73.26% | 66.09% | 77.56% | 67.69% |
| mPLUG-Owl2 (LLaMA-7B) [59] | 72.26% | 55.53% | 58.64% | 52.59% | 71.36% | 58.90% | 73.00% | 62.68% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 64.60% | 59.22% | 55.76% | 47.98% | 67.30% | 58.90% | 73.76% | 60.07% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 64.96% | 64.86% | 54.12% | 53.55% | 66.59% | 58.90% | 71.48% | 61.40% |
| InternLM-XComposer-VL (InternLM) [61] | 68.43% | 62.04% | 61.93% | 56.81% | 70.41% | 57.53% | 77.19% | 64.35% |
| IDEFICS-Instruct (LLaMA-7B) [62] | 60.04% | 46.42% | 46.71% | 40.38% | 59.90% | 47.26% | 64.77% | 51.51% |
| Qwen-VL (QwenLM) [63] | 65.33% | 60.74% | 58.44% | 54.13% | 66.35% | 58.22% | 73.00% | 61.67% |
| Shikra (Vicuna-7B) [64] | 69.09% | 47.93% | 46.71% | 47.31% | 60.86% | 53.08% | 64.77% | 55.32% |
| Otter-v1 (MPT-7B) [6] | 57.66% | 39.70% | 42.59% | 42.12% | 48.93% | 47.60% | 54.17% | 47.22% |
| InstructBLIP (Flan-T5-XL) [5] | 69.53% | 59.00% | 56.17% | 57.31% | 65.63% | 56.51% | 71.21% | 61.94% |
| InstructBLIP (Vicuna-7B) [5] | 70.99% | 51.41% | 43.00% | 45.00% | 63.01% | 57.19% | 64.39% | 55.85% |
| VisualGLM-6B (GLM-6B) [65] | 61.31% | 53.58% | 44.03% | 48.56% | 54.89% | 55.48% | 57.79% | 53.31% |
| mPLUG-Owl (LLaMA-7B) [52] | 72.45% | 54.88% | 47.53% | 49.62% | 63.01% | 62.67% | 66.67% | 58.93% |
| LLaMA-Adapter-V2 [66] | 66.61% | 54.66% | 51.65% | 56.15% | 61.81% | 59.25% | 54.55% | 58.06% |
| LLaVA-v1 (Vicuna-13B) [3] | 57.12% | 54.88% | 51.85% | 45.58% | 58.00% | 57.19% | 64.77% | 54.72% |
| MiniGPT-4 (Vicuna-13B) [4] | 60.77% | 50.33% | 43.00% | 45.58% | 52.51% | 53.42% | 60.98% | 51.77% |
@@
| Qwen-VL-Max (Close-Source) [63] | 73.20% | 81.02% | 68.39% | 70.84% | 74.57% | 73.11% | 80.44% | 73.90% |
| Gemini-Pro (Close-Source) [27] | 71.26% | 71.39% | 65.59% | 67.30% | 73.04% | 65.88% | 73.60% | 69.46% |
| GPT-4V (Close-Source) [28] | 77.72% | 78.39% | 66.45% | 71.01% | 71.07% | 79.36% | 78.91% | 74.10% |
@@ II-D1 Weak Measurability of MLLM Outputs
In Q-Bench, we aim to fairly compare the assessment ability between different MLLMs on diverse low-level appearances. Henceforth, our principle is to define a unified, simplest instruction that is applicable for all MLLMs on all IQA datasets. Under this principle, we conduct toy experiments on Shikra [64] and LLaVA-v1 [3], with two simple instruction strategies: (A) Direct Instruction, in which the prompt is designed as simple as “Rate the quality of the image”. The top-frequency answers are good (78%), and poor (20%), with other outputs almost negligible. (B) Numerical Instruction, in which we specifically instruct numerical ratings, with the prompt: “Score the quality of the image from 1 to 5, with 1 as lowest and 5 as highest.”. Under the numerical strategy, the top-frequency answers are 5 (84%), 1 (9%), and 3 (5%); though within the score range, the frequencies of scores 2 and 4 are both less than 1%. The toy experiments imply the weak measurability of MLLM outputs, given that the answers are statistically 1) biased towards positive, 2) biased towards extreme, and 3) with only two effective scales. Therefore, it is necessary to explore extended strategies for MLLMs to provide truly quantifiable outputs for low-level assessment.
@@ II-D2 A Softmax-based Evaluation Strategy
This simple and generally-applicable strategy enables us to collect quantifiable outputs () from MLLMs with higher correlation to human ratings, as verified in our experimental analysis (Table VII).
@@ II-D3 Prompt Ensemble for Boosting Quantitative Abilities of MLLMs
where indicates the positive token set (from good, fine, high, etc.), while represents the negative token set (from poor, bad, low, etc.). The implementation of the prompt ensemble approach does not add extra computational complexity. The core computation occurs once the input prompt is entered and the language model generates the [SCORE_TOKEN]. After this, we only require tokenization of the words used, followed by the calculation of logits for the [SCORE_TOKEN]. The boosted performance is listed in Table VIII.
@@
| Emu2-Chat (LLaMA-33B) [25] | 56.64% | 41.15% | 49.62% | 49.12% | 51.91% | 47.86% | 60.00% | 50.05% |
| Fuyu-8B (Persimmon-8B) [57] | 68.76% | 33.56% | 38.78% | 46.83% | 54.03% | 47.86% | 55.00% | 49.15% |
| BakLLava (Mistral-7B) [26] | 56.92% | 43.83% | 50.00% | 49.33% | 54.34% | 50.66% | 52.22% | 50.94% |
| mPLUG-Owl2 (Q-Instruct) [67] | 59.19% | 42.12% | 47.43% | 49.63% | 52.48% | 49.81% | 53.88% | 50.54% |
| mPLUG-Owl2 (LLaMA-7B) [59] | 58.43% | 39.72% | 48.39% | 49.04% | 51.55% | 47.50% | 60.55% | 49.85% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 60.46% | 42.85% | 41.53% | 47.88% | 51.89% | 46.55% | 59.57% | 49.32% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 56.42% | 42.46% | 48.38% | 48.15% | 53.41% | 48.84% | 54.44% | 49.85% |
@@
| Qwen-VL-Max (Close-Source) [63] | 71.96% | 62.87% | 65.53% | 69.21% | 62.69% | 67.54% | 66.01% | 67.27% |
| Gemini-Pro (Close-Source) [27] | 64.98% | 51.36% | 54.16% | 58.17% | 56.52% | 57.73% | 57.22% | 57.64% |
| GPT-4V (Close-Source) [28] | 79.34% | 70.54% | 78.52% | 75.84% | 77.95% | 78.80% | 66.11% | 76.52% |
@@
| Emu2-Chat (LLaMA-33B) [25] | 51.94% | 29.78% | 53.84% | 42.01% | 55.71% | 46.26% | 49.09% | 47.08% |
| Fuyu-8B (Persimmon-8B) [57] | 70.36% | 28.13% | 35.98% | 44.08% | 57.43% | 47.02% | 51.11% | 47.94% |
| BakLLava (Mistral-7B) [26] | 60.09% | 45.42% | 50.86% | 53.09% | 58.82% | 54.52% | 55.55% | 52.75% |
| mPLUG-Owl2 (Q-Instruct) [67] | 60.24% | 47.46% | 48.78% | 52.81% | 53.97% | 51.42% | 59.11% | 53.15% |
| mPLUG-Owl2 (LLaMA-7B) [59] | 58.07% | 36.61% | 48.44% | 47.74% | 51.90% | 45.73% | 60.00% | 48.94% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 60.72% | 42.37% | 50.17% | 49.15% | 59.86% | 52.97% | 49.77% | 52.25% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 57.34% | 47.45% | 49.13% | 49.01% | 59.51% | 52.06% | 52.00% | 52.05% |
@@
| Qwen-VL-Max (Close-Source) [63] | 67.65% | 67.56% | 65.35% | 69.09% | 61.18% | 68.65% | 61.29% | 66.99% |
| Gemini-Pro (Close-Source) [27] | 65.78% | 56.61% | 56.74% | 60.42% | 60.55% | 60.46% | 60.44% | 60.46% |
| GPT-4V (Close-Source) [28] | 79.75% | 69.49% | 84.42% | 77.32% | 79.93% | 81.00% | 68.00% | 78.07% |

### v2

@@ II-C Benchmark on Low-level Description Ability
In the second task of Q-Bench+, we evaluate the language description ability of MLLMs on low-level information. This task is a sibling task of image captioning [cococaps, flickrcaps, nocaps] that describes image content with natural language, with a specific concern on the low-level appearance of images. To evaluate this ability automatically, we first derive a golden low-level description dataset, denoted as LLDescribe+ (Sec. II-C1), including one long (average 58 words) golden description provided by experts for each of 499 images. With these golden text descriptions, we are able to measure the quality of output low-level descriptions from MLLMs with a single-modal GPT, under the three dimensions: completeness, preciseness, as well as relevance (Sec II-C2). The discussions of the golden descriptions and the evaluation process are as follows.
@@ II-C1 Defining Golden Low-level Descriptions for Images
For the description ability, MLLMs should accurately and completely describe low-level visual information of images. Thus, the ground truths for these MLLMs are also built within a basic principle to cover as many low-level concerns as possible, so long as they are enumerated in Sec. 3 and occur in images. The resulting golden descriptions in LLDescribe+ have an average duration of 58 words, notably longer than common high-level image caption datasets (11 for [nocaps], 10 for [cococaps]). Similar to the LLVisionQA+ dataset for the perception task, the 499 single images and 450 image pairs in LLDescribe+ dataset also include all 10 sources (as in Table I) to cover images with diverse low-level appearances. The golden descriptions on different sources of images are depicted in Fig. 4.
@@ II-C2 Evaluation with Single-modal GPT
Recent studies [vicuna] have proved single-modal GPT [openai2023gpt4] to be a reliable evaluation tool for pure language tasks. Via the LLDescribe+ dataset, we convert the multi-modality problem into a text-only setting, by matching the MLLM outputs with the golden descriptions with single-modal GPT under three dimensions: (1) Completeness. More matched information with the golden description is encouraged. (2) Preciseness. The controversial information with the golden description is punished. (3) Relevance. More proportions of MLLM outputs should be related to low-level information, instead of others. Each dimension is scored among [0,1,2]. Similar as Sec. 4, we repeat 5 rounds for each single evaluation and collect the weighted average as the final score.
@@ II-D Benchmark on Precise Quality Assessment Ability
| Emu2-Chat (LLaMA-33B) [emu2] | 71.81% | 67.25% | 56.18% | 64.78% | 63.19% | 63.48% | 72.24% | 65.28% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 53.33% | 43.70% | 38.00% | 40.81% | 47.40% | 45.45% | 49.23% | 45.05% |
| BakLLava (Mistral-7B) [bakllava] | 66.00% | 56.16% | 51.12% | 51.15% | 61.57% | 53.72% | 72.00% | 57.48% |
| SPHINX [sphinx] | 74.18% | 68.81% | 62.07% | 63.62% | 71.76% | 66.12% | 76.33% | 68.56% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 72.18% | 57.96% | 56.19% | 56.68% | 69.21% | 53.29% | 72.65% | 61.61% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 66.36% | 58.19% | 50.51% | 49.42% | 65.74% | 54.61% | 70.61% | 58.66% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 65.27% | 64.38% | 56.59% | 56.03% | 67.13% | 61.18% | 67.35% | 62.14% |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 69.45% | 65.27% | 60.85% | 61.67% | 70.14% | 56.91% | 75.10% | 65.35% |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 56.18% | 44.69% | 44.02% | 42.80% | 54.17% | 44.74% | 56.33% | 48.70% |
| Qwen-VL (QwenLM) [Qwen-VL] | 63.09% | 58.19% | 56.39% | 50.58% | 62.73% | 57.89% | 73.88% | 59.40% |
| Shikra (Vicuna-7B) [shikra] | 65.64% | 47.35% | 49.09% | 48.83% | 59.49% | 50.00% | 64.08% | 54.65% |
| Otter-v1 (MPT-7B) [otter] | 57.09% | 40.71% | 39.55% | 42.22% | 49.31% | 44.08% | 52.65% | 46.35% |
| InstructBLIP (Flan-T5-XL) [iblip] | 67.64% | 59.96% | 55.98% | 56.23% | 65.51% | 58.22% | 69.39% | 61.47% |
| InstructBLIP (Vicuna-7B) [iblip] | 71.64% | 52.65% | 43.81% | 48.64% | 62.50% | 55.59% | 64.90% | 56.72% |
| VisualGLM-6B (GLM-6B) [glm] | 60.18% | 54.20% | 46.25% | 51.75% | 54.40% | 53.62% | 57.14% | 53.78% |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 66.00% | 54.87% | 44.02% | 51.36% | 55.09% | 54.28% | 65.71% | 55.38% |
| LLaMA-Adapter-V2 [llamaadapterv2] | 66.18% | 59.29% | 52.13% | 57.39% | 56.25% | 63.16% | 64.90% | 59.46% |
| LLaVA-v1 (Vicuna-13B) [llava] | 54.00% | 53.10% | 55.38% | 48.64% | 54.63% | 55.59% | 63.27% | 54.18% |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 55.82% | 50.22% | 40.37% | 42.02% | 48.38% | 51.97% | 61.22% | 49.03% |
@@
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 75.60% | 79.43% | 66.09% | 73.39% | 74.08% | 71.0% | 76.92% | 73.63% |
| Gemini-Pro (Close-Source) [geminipro] | 68.80% | 73.74% | 62.34% | 66.30% | 71.34% | 63.91% | 73.09% | 68.16% |
| GPT-4V (Close-Source) [openai2023gpt4] | 76.85% | 79.17% | 67.52% | 73.53% | 76.18% | 72.83% | 76.47% | 74.51% |
@@
| Emu2-Chat (LLaMA-33B) [emu2] | 70.09% | 65.12% | 54.11% | 66.22% | 62.96% | 63.47% | 73.21% | 64.32% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 62.22% | 35.79% | 36.62% | 41.07% | 49.40% | 45.89% | 49.04% | 45.75% |
| BakLLava (Mistral-7B) [bakllava] | 66.46% | 61.48% | 54.83% | 51.33% | 63.76% | 56.52% | 78.16% | 61.02% |
| SPHINX [sphinx] | 74.45% | 65.50% | 62.13% | 59.11% | 73.26% | 66.09% | 77.56% | 67.69% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 72.26% | 55.53% | 58.64% | 52.59% | 71.36% | 58.90% | 73.00% | 62.68% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 64.60% | 59.22% | 55.76% | 47.98% | 67.30% | 58.90% | 73.76% | 60.07% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 64.96% | 64.86% | 54.12% | 53.55% | 66.59% | 58.90% | 71.48% | 61.40% |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 68.43% | 62.04% | 61.93% | 56.81% | 70.41% | 57.53% | 77.19% | 64.35% |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 60.04% | 46.42% | 46.71% | 40.38% | 59.90% | 47.26% | 64.77% | 51.51% |
| Qwen-VL (QwenLM) [Qwen-VL] | 65.33% | 60.74% | 58.44% | 54.13% | 66.35% | 58.22% | 73.00% | 61.67% |
| Shikra (Vicuna-7B) [shikra] | 69.09% | 47.93% | 46.71% | 47.31% | 60.86% | 53.08% | 64.77% | 55.32% |
| Otter-v1 (MPT-7B) [otter] | 57.66% | 39.70% | 42.59% | 42.12% | 48.93% | 47.60% | 54.17% | 47.22% |
| InstructBLIP (Flan-T5-XL) [iblip] | 69.53% | 59.00% | 56.17% | 57.31% | 65.63% | 56.51% | 71.21% | 61.94% |
| InstructBLIP (Vicuna-7B) [iblip] | 70.99% | 51.41% | 43.00% | 45.00% | 63.01% | 57.19% | 64.39% | 55.85% |
| VisualGLM-6B (GLM-6B) [glm] | 61.31% | 53.58% | 44.03% | 48.56% | 54.89% | 55.48% | 57.79% | 53.31% |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 72.45% | 54.88% | 47.53% | 49.62% | 63.01% | 62.67% | 66.67% | 58.93% |
| LLaMA-Adapter-V2 [llamaadapterv2] | 66.61% | 54.66% | 51.65% | 56.15% | 61.81% | 59.25% | 54.55% | 58.06% |
| LLaVA-v1 (Vicuna-13B) [llava] | 57.12% | 54.88% | 51.85% | 45.58% | 58.00% | 57.19% | 64.77% | 54.72% |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 60.77% | 50.33% | 43.00% | 45.58% | 52.51% | 53.42% | 60.98% | 51.77% |
@@
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 73.20% | 81.02% | 68.39% | 70.84% | 74.57% | 73.11% | 80.44% | 73.90% |
| Gemini-Pro (Close-Source) [geminipro] | 71.26% | 71.39% | 65.59% | 67.30% | 73.04% | 65.88% | 73.60% | 69.46% |
| GPT-4V (Close-Source) [openai2023gpt4] | 77.72% | 78.39% | 66.45% | 71.01% | 71.07% | 79.36% | 78.91% | 74.10% |
@@ II-D1 Weak Measurability of MLLM Outputs
In Q-Bench+, we aim to fairly compare the assessment ability between different MLLMs on diverse low-level appearances. Henceforth, our principle is to define a unified, simplest instruction that is applicable for all MLLMs on all IQA datasets. Under this principle, we conduct toy experiments on Shikra [shikra] and LLaVA-v1 [llava], with two simple instruction strategies: (A) Direct Instruction, in which the prompt is designed as simple as “Rate the quality of the image”. The top-frequency answers are good (78%), and poor (20%), with other outputs almost negligible. (B) Numerical Instruction, in which we specifically instruct numerical ratings, with the prompt: “Score the quality of the image from 1 to 5, with 1 as lowest and 5 as highest.”. Under the numerical strategy, the top-frequency answers are 5 (84%), 1 (9%), and 3 (5%); though within the score range, the frequencies of scores 2 and 4 are both less than 1%. The toy experiments imply the weak measurability of MLLM outputs, given that the answers are statistically 1) biased towards positive, 2) biased towards extreme, and 3) with only two effective scales. Therefore, it is necessary to explore extended strategies for MLLMs to provide truly quantifiable outputs for low-level assessment.
@@ II-D2 A Softmax-based Evaluation Strategy
This simple and generally-applicable strategy enables us to collect quantifiable outputs () from MLLMs with higher correlation to human ratings, as verified in our experimental analysis (Table XII).
@@ II-D3 Prompt Ensemble for Boosting Quantitative Abilities of MLLMs
where indicates the positive token set (from good, fine, high, etc.), while represents the negative token set (from poor, bad, low, etc.). The implementation of the prompt ensemble approach does not add extra computational complexity. The core computation occurs once the input prompt is entered and the language model generates the [SCORE_TOKEN]. After this, we only require tokenization of the words used, followed by the calculation of logits for the [SCORE_TOKEN]. The boosted performance is listed in Table XIII.
@@
| Emu2-Chat (LLaMA-33B) [emu2] | 56.64% | 41.15% | 49.62% | 49.12% | 51.91% | 47.86% | 60.00% | 50.05% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 68.76% | 33.56% | 38.78% | 46.83% | 54.03% | 47.86% | 55.00% | 49.15% |
| BakLLava (Mistral-7B) [bakllava] | 56.92% | 43.83% | 50.00% | 49.33% | 54.34% | 50.66% | 52.22% | 50.94% |
| mPLUG-Owl2 (Q-Instruct) [q-instruct] | 59.19% | 42.12% | 47.43% | 49.63% | 52.48% | 49.81% | 53.88% | 50.54% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 58.43% | 39.72% | 48.39% | 49.04% | 51.55% | 47.50% | 60.55% | 49.85% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.46% | 42.85% | 41.53% | 47.88% | 51.89% | 46.55% | 59.57% | 49.32% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 56.42% | 42.46% | 48.38% | 48.15% | 53.41% | 48.84% | 54.44% | 49.85% |
@@
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 71.96% | 62.87% | 65.53% | 69.21% | 62.69% | 67.54% | 66.01% | 67.27% |
| Gemini-Pro (Close-Source) [geminipro] | 64.98% | 51.36% | 54.16% | 58.17% | 56.52% | 57.73% | 57.22% | 57.64% |
| GPT-4V (Close-Source) [openai2023gpt4] | 79.34% | 70.54% | 78.52% | 75.84% | 77.95% | 78.80% | 66.11% | 76.52% |
@@
| Emu2-Chat (LLaMA-33B) [emu2] | 51.94% | 29.78% | 53.84% | 42.01% | 55.71% | 46.26% | 49.09% | 47.08% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 70.36% | 28.13% | 35.98% | 44.08% | 57.43% | 47.02% | 51.11% | 47.94% |
| BakLLava (Mistral-7B) [bakllava] | 60.09% | 45.42% | 50.86% | 53.09% | 58.82% | 54.52% | 55.55% | 52.75% |
| mPLUG-Owl2 (Q-Instruct) [q-instruct] | 60.24% | 47.46% | 48.78% | 52.81% | 53.97% | 51.42% | 59.11% | 53.15% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 58.07% | 36.61% | 48.44% | 47.74% | 51.90% | 45.73% | 60.00% | 48.94% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.72% | 42.37% | 50.17% | 49.15% | 59.86% | 52.97% | 49.77% | 52.25% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 57.34% | 47.45% | 49.13% | 49.01% | 59.51% | 52.06% | 52.00% | 52.05% |
@@
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 67.65% | 67.56% | 65.35% | 69.09% | 61.18% | 68.65% | 61.29% | 66.99% |
| Gemini-Pro (Close-Source) [geminipro] | 65.78% | 56.61% | 56.74% | 60.42% | 60.55% | 60.46% | 60.44% | 60.46% |
| GPT-4V (Close-Source) [openai2023gpt4] | 79.75% | 69.49% | 84.42% | 77.32% | 79.93% | 81.00% | 68.00% | 78.07% |